### v1

@@ 3.2 Reconstructing sea surface wave height fields with stereo data
First, we train a functional autoencoder to reconstruct the principal structures of sea surface wave fields using 7000 image frames, with the remaining 1000 frames held out as a test set. During masked training, $50\%$ of the data points in each image are provided to the encoder as input for feature extraction, while the model is trained to reconstruct the complementary $50\%$ of the points. To evaluate the reconstruction performance, we investigate the sensitivity of the model to different input point ratios. Specifically, we fix an arbitrary sample from the held-out set and generate 100 distinct masks with point ratios of $50\%,10\%,3\%,1\%$, and $0.5\%$. Each masked input is then encoded and decoded on the full grid, and kernel density estimates of the reconstruction root mean square error (RMSE) are computed, as presented in Fig. 7.
@@
From Fig. 7, it is evident that when the input ratio is above approximately $3\%$, the functional autoencoder achieves highly accurate and stable reconstructions of the principal structures, with RMSE means remaining close to 0.40 and variances below 0.015. As the input ratio decreases further, the reconstruction quality gradually degrades. At $1\%$ input, for example, the mean RMSE increases to about 0.43 and the variance nearly doubles, and at $0.5\%$ input the mean RMSE reaches approximately 0.45 with the largest spread in the error distribution. Despite this deterioration, the reconstructions still preserve the correct large-scale patterns of the wave field, highlighting the strong robustness of the functional autoencoder to severe data sparsity in capturing the dominant structures.
@@
Figure 7: The reconstructed principal structures and kernel density estimates for RMSE on the reference sample across 100 randomly chosen meshes.
@@
When the functional autoencoder is trained and fixed, it is used to output principal structures according to sparse measurements as conditions to train a conditional diffusion model with the proposed mask-cascade training strategy. The input point ratio of the fixed functional autoencoder is set as $0.5\%$ to enhance the robustness of conditional diffusion model to sparsity. To test the generation results, given $0.5\%$ randomly selected observation points from each test sample, the functional autoencoder reconstructs the corresponding principal structures. These structures are then used as conditions to input the conditional diffusion model while manifold constrained gradients are incorporated into the sampling process to ensure consistency between the generated full-field data and the observed $0.5\%$ points. As shown in Fig. 8(a), the generated fields exhibit high fidelity to the ground truth. It is important to note that each generated realization corresponds to a single sample from $p(\boldsymbol{u}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$. To quantitatively evaluate the uncertainty of the generative model, we further fix a test sample and generate random masks at different observation ratios ($0.2\%,0.5\%,1\%,3\%$, and $5\%$). After fixing these masks, we perform 100 independent generations, and compute the kernel density estimates of the resulting RMSE distributions. The results, shown in Fig. 8, where the mean field of generations match the ground truth very well. As the percentage of observed points increases from $0.2\%$ to $1\%$, the mean RMSE decreases notably, indicating improved reconstruction quality with denser observations. In particular, the drop in both mean error and variance around the $1\%$ input level suggests that even a relatively small increase in observations can significantly stabilize the generation. When the ratio is further increased to $3\%$ and $5\%$, the mean RMSE continues to decrease slightly, approaching a lower bound, while the variance remains at a low level since the principal structures reconstruction results are more stable under these ratios, reflecting both accuracy and robustness. It is worth noting that although the conditional diffusion model is trained with $0.5\%$ input points, it generalizes well to different input ratios and consistently delivers satisfactory reconstructions even with only $0.2\%$ points. This can be attributed to two factors. First, during mask-cascade training, the conditions are generated by random sampling, which enhances the robustness of the conditional diffusion model to diverse and sparse conditions. Second, the manifold-constrained gradient further enforces consistency between the generated fields and the sparse observations while keeping the sampling trajectory on the data manifold, which ensures reliable reconstruction of the physical fields even under varying levels of sparsity.
@@ 3.3 Reconstructing global sea surface temperature fields with reanalysis data
We begin by training a functional autoencoder on the dataset to learn a compact representation of the global sea surface temperature field. During the masked training, $50\%$ of the data points in each frame are randomly selected as inputs to the encoder, while the remaining $50\%$ are used as targets for the decoder. As illustrated in Fig. 10(a) (Condition), the trained functional autoencoder is capable of accurately reconstructing the large-scale background patterns from only partial observations, capturing the essential spatial variability of the ocean surface temperature fields. Once the autoencoder is fully trained, it is frozen and integrated into the Cascaded Sensing framework to provide real-time background field reconstructions that serve as conditioning information for the conditional diffusion model during mask-cascade training. Specifically, a sparse subset of points ($0.5\%$) is randomly sampled at each step during training, from which the functional autoencoder reconstructs a smooth background field. The conditional diffusion model then learns to refine this background and recover the full field. After 800 epochs of training, the conditional diffusion model is well-trained to reconstruct fine-scale details based on conditions and extremely sparse measurements.
To evaluate the model’s generalization ability under more challenging observation conditions, we randomly sample only $0.1\%$ of the data points from the 15 predefined global subregions, a sampling ratio that is five times lower than that used during training. A representative reconstruction (RMSE = 0.0144) is shown in Fig. 10(a). Remarkably, as evident in the zoomed-in views, the Cascaded Sensing framework not only reconstructs the global background temperature field with high accuracy but also successfully captures fine-scale turbulent structures, which are typically difficult to recover from such sparse observations.
@@
4 Discussion
In this study, we conduct a comprehensive evaluation of the proposed Cascaded Sensing (Cas-Sensing) framework under a broad range of scenarios characterized by different sensor configurations, sampling sparsities, and geometric boundaries. Our results clearly show that the hierarchical structure of Cas-Sensing can substantially mitigate the ill-posedness of reconstructing multi-scale physical fields from extremely sparse measurements. By decomposing the reconstruction process into two stages—large-scale structure recovery via a functional autoencoder and fine-scale detail generation through a conditional diffusion model—the proposed method achieves accurate and stable reconstructions with high data fidelity.
A central strength of this cascaded design lies in its ability to reduce the reconstruction ambiguity caused by sparse and irregular inputs. By first establishing reliable coarse-scale conditioning, the diffusion model no longer needs to infer the entire field from incomplete observations, but focuses on capturing the missing small-scale structures. This hierarchical formulation enhances reconstruction stability and reduces the variability across different sampling patterns. Moreover, the mask-cascade training strategy significantly improves the model’s robustness and generalization, enabling it to adapt to diverse sensor layouts without the need for retraining.
@@ 4 Discussion
These advantages make Cas-Sensing particularly effective for scientific applications. For example, in fluid dynamics, accurate flow field reconstruction from sparse sensors is crucial for monitoring and control; in climate and ocean science, reconstructing global or regional temperature fields from sparse measurements enables more reliable environmental assessment; and in geophysical applications, the ability to generalize across sampling configurations supports practical deployment in complex field settings.
Overall, Cas-Sensing provides a flexible and efficient framework for reconstructing multi-scale fields. By leveraging hierarchical representation learning and generative modeling, it delivers both accuracy and robustness across a variety of reconstruction scenarios, offering a promising pathway for advancing data-driven scientific sensing and prediction.
5 Conclusion and future work
In this study, we have presented Cascade Sensing (Cas-Sensing) as an effective framework for reconstructing full-field data from highly sparse observations. By leveraging an autoencoder-diffusion cascade framework, Cas-Sensing is capable of producing accurate and coherent reconstructions even when the input measurements are extremely limited. The mask-cascade training strategy, combined with manifold constrained gradient, enhances the robustness of the model to diverse sparse conditions and ensures the consistency of the reconstructed physical fields. These results demonstrate the potential of Cas-Sensing as a powerful tool for efficient, data-driven sensing and predictive modeling in complex physical systems.
Nevertheless, several limitations remain in the current implementation. Firstly, the present study focuses exclusively on two-dimensional spatial domains, leaving the extension to higher-dimensional settings unexplored. Scaling Cas-Sensing to three-dimensional or higher-dimensional problems may introduce additional computational and modeling challenges that warrant careful investigation. Secondly, the current framework addresses only spatial reconstruction, without explicitly modeling temporal dynamics. Consequently, the method cannot fully capture spatiotemporal evolution in dynamic systems, which may limit its applicability in real-time monitoring or predictive scenarios. Thirdly, the generative component currently relies on a vanilla DDPM, which, while effective, may impose constraints on sampling speed and computational efficiency. Alternative generative models, such as flow matching, could potentially accelerate sampling and further improve reconstruction fidelity.
Future research should aim to overcome these limitations. Extending Cas-Sensing to higher-dimensional domains would broaden its applicability to more complex physical systems, including 3D flow fields or volumetric sensor networks. Integrating temporal modeling within the cascade framework could enable spatiotemporal reconstruction, capturing dynamic evolution in addition to spatial correlations, and allowing the framework to support real-time prediction tasks. Moreover, the adoption of more advanced generative models could enhance sampling efficiency without sacrificing accuracy, making Cas-Sensing more practical for large-scale deployments. Collectively, these developments would not only enhance the versatility and scalability of Cas-Sensing but also reinforce its potential as a robust and reliable tool for data-driven sensing, reconstruction, and prediction in complex physical environments.

### v2

@@ 3.2 Reconstructing sea surface wave height fields with stereo data
First, we train a functional autoencoder to reconstruct the coarse-scale structures of sea surface wave fields using 7000 image frames, with the remaining 1000 frames held out as a test set. During masked training, $50\%$ of the data points in each image are provided to the encoder as input for feature extraction, while the model is trained to reconstruct the complementary $50\%$ of the points. To evaluate the reconstruction performance, we investigate the sensitivity of the model to different input point ratios. Specifically, we fix an arbitrary sample from the held-out set and generate 100 distinct masks with point ratios of $50\%,10\%,3\%,1\%$, and $0.5\%$. Each masked input is then encoded and decoded on the full grid, and kernel density estimates of the reconstruction root mean square error (RMSE) are computed, as presented in Fig. 10.
@@
From Fig. 10, it is evident that when the input ratio is above approximately $3\%$, the functional autoencoder achieves highly accurate and stable reconstructions of the coarse-scale structures, with RMSE means remaining close to 0.40 and variances below 0.015. As the input ratio decreases further, the reconstruction quality gradually degrades. At $1\%$ input, for example, the mean RMSE increases to about 0.43 and the variance nearly doubles, and at $0.5\%$ input the mean RMSE reaches approximately 0.45 with the largest spread in the error distribution. Despite this deterioration, the reconstructions still preserve the correct coarse-scale patterns of the wave field, highlighting the strong robustness of the functional autoencoder to severe data sparsity in capturing the dominant structures.
@@
Figure 10: The reconstructed coarse-scale structures and kernel density estimates for RMSE on the reference sample across 100 randomly chosen meshes.
When the functional autoencoder is trained and fixed, it is used to output coarse-scale structures according to sparse measurements as conditions to train a conditional diffusion model with the proposed mask-cascade training strategy.
The input point ratio of the fixed functional autoencoder is set as $0.5\%$ to enhance the robustness of conditional diffusion model to sparsity. To test the generation results, given $0.5\%$ randomly selected observation points from each test sample, the functional autoencoder reconstructs the corresponding coarse-scale structures. These structures are then used as conditions to input the conditional diffusion model, while manifold constrained gradients are incorporated into the sampling process to ensure consistency between the generated full-field data and the observed $0.5\%$ points.
@@
As shown in Fig. 11(a), the generated fields exhibit high fidelity to the ground truth. It is important to note that each generated realization corresponds to a single sample from $p(\boldsymbol{u}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$. To quantitatively evaluate the uncertainty of the generative model, we further fix a test sample and generate random masks at different observation ratios ($0.2\%,0.5\%,1\%,3\%$, and $5\%$). After fixing these masks, we perform 100 independent generations, and compute the kernel density estimates of the resulting RMSE distributions. The results, shown in Fig. 11, where the mean field of generations match the ground truth very well. As the percentage of observed points increases from $0.2\%$ to $1\%$, the mean RMSE decreases notably, indicating improved reconstruction quality with denser observations. In particular, the drop in both mean error and variance around the $1\%$ input level suggests that even a relatively small increase in observations can significantly stabilize the generation. When the ratio is further increased to $3\%$ and $5\%$, the mean RMSE continues to decrease slightly, approaching a lower bound, while the variance remains at a low level since the coarse-scale structures reconstruction results are more stable under these ratios, reflecting both accuracy and robustness.
It is worth noting that although the conditional diffusion model is trained with $0.5\%$ input points, it generalizes well to different input ratios and consistently delivers satisfactory reconstructions even with only $0.2\%$ points. This can be attributed to two factors. First, during mask-cascade training, the conditions are generated by random sampling, which enhances the robustness of the conditional diffusion model to diverse and sparse conditions. Second, the manifold-constrained gradient further enforces consistency between the generated fields and the sparse observations while keeping the sampling trajectory on the data manifold, which ensures reliable reconstruction of the physical fields even under varying levels of sparsity.
@@ 3.3 Reconstructing global sea surface temperature fields with reanalysis data
We begin by training a functional autoencoder on the dataset to learn a compact representation of the global sea surface temperature field. During the masked training, $50\%$ of the data points in each frame are randomly selected as inputs to the encoder, while the remaining $50\%$ are used as targets for the decoder. As illustrated in Fig. 13(a) (Condition), the trained functional autoencoder is capable of accurately reconstructing the coarse-scale background patterns from only partial observations, capturing the essential spatial variability of the ocean surface temperature fields. Once the autoencoder is fully trained, it is frozen and integrated into the Cascaded Sensing framework to provide real-time background field reconstructions that serve as conditioning information for the conditional diffusion model during mask-cascade training. Specifically, a sparse subset of points ($0.5\%$) is randomly sampled at each step during training, from which the functional autoencoder reconstructs a smooth background field. The conditional diffusion model then learns to refine this background and recover the full field. After 800 epochs of training, the conditional diffusion model is well-trained to reconstruct fine-scale details based on conditions and extremely sparse measurements.
To evaluate the model’s generalization ability under more challenging observation conditions, we randomly sample only $0.1\%$ of the data points from the 15 predefined global subregions, a sampling ratio that is five times lower than that used during training. A representative reconstruction (RMSE = 0.0144) is shown in Fig. 13(a). Remarkably, as evident in the zoomed-in views, the Cascaded Sensing framework not only reconstructs the global background temperature field with high accuracy but also successfully captures fine-scale turbulent structures, which are typically highly challenging to recover from such sparse observations.
@@
4 Discussion and conclusion
In this study, we presented and systematically evaluated Cascaded Sensing (Cas-Sensing) as a probabilistic paradigm for reconstructing multi-scale physical fields under extreme data sparsity. By explicitly recognizing sparse sensing as a severely ill-posed inverse problem with intrinsic non-uniqueness, Cas-Sensing reformulates full-field reconstruction as a hierarchical probabilistic inference task, rather than attempting to learn a direct deterministic mapping from sparse observations to full fields. Extensive experiments across a wide range of sensor configurations, sampling sparsities, and geometric boundaries consistently demonstrate that this hierarchical formulation effectively mitigates the ill-posedness inherent in sparse reconstruction problems.
A central strength of the proposed framework lies in the introduction of an explicit intermediate representation that captures coarse-scale structures. By first recovering physically meaningful coarse-scale components and geometric features via a functional autoencoder, the original ill-posed problem is decomposed into two substantially better-conditioned subproblems. This decomposition constrains subsequent reconstruction to remain consistent with dominant structures, allowing the conditional diffusion model to focus exclusively on generating missing refined-scale details as residuals. As a result, reconstruction responsibilities are decoupled across spatial scales, improving stability and fidelity while avoiding information loss associated with unified latent-space generation.
The robustness and generalization capabilities of Cas-Sensing are further enhanced by the proposed mask-cascade training strategy and manifold-constrained gradient-based sampling. By exposing the diffusion model to a distribution of imperfect conditioning structures induced by varying sparsity patterns, the model learns a conditional distribution that remains stable across diverse sensor layouts and extreme sparsity levels. During inference, measurement consistency is enforced through posterior-guided sampling without retraining, preserving coherence with the learned data manifold while reducing reconstruction ambiguity and variability.
@@ 4 Discussion and conclusion
These properties make Cas-Sensing particularly suitable for practical scientific and engineering applications. In fluid dynamics, reliable reconstruction of coherent flow structures from sparse sensors supports monitoring and control tasks. In climate and ocean science, the ability to recover coarse-scale background fields together with refined-scale variability enables robust environmental assessment. In geophysical and engineering sensing, generalization across sensor layouts and geometric configurations facilitates deployment in complex, real-world environments.
Despite these advantages, several limitations remain. The current implementation is restricted to two-dimensional spatial domains and does not explicitly incorporate temporal dynamics, limiting applicability to fully spatiotemporal or three-dimensional systems. In addition, reliance on a vanilla denoising diffusion probabilistic model introduces practical constraints on sampling efficiency, which may become significant in coarse-scale applications.
These limitations also suggest promising directions for future research. Extending the cascaded sensing paradigm to higher-dimensional settings would enable reconstruction of volumetric fields and complex 3D geometries. Incorporating temporal modeling into the cascade could further generalize the framework to spatiotemporal reconstruction and prediction, supporting real-time monitoring and forecasting tasks. Moreover, integrating more advanced generative paradigms, such as flow-matching-based or accelerated diffusion methods, may substantially improve sampling efficiency while preserving reconstruction quality.
Overall, Cas-Sensing represents a general, extensible, and principled framework for data-driven scientific sensing. By explicitly structuring uncertainty, separating coarse- and fine-scale reconstruction responsibilities, and modeling conditional distributions rather than deterministic mappings, this work provides a systematic approach to addressing non-uniqueness and multi-scale complexity in sparse inverse problems. The proposed paradigm lays the groundwork for future sensing and inference systems that are both robust to severe data incompleteness and adaptable to complex physical environments.
Acknowledgements
The authors wish to express their gratitude for the financial support received from Guangdong Provincial Fund - Special Innovation Project (2024KTSCX038); Research Grants Council of Hong Kong through the Research Impact Fund (R5006-23); the Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No. 2023B1212010007). The authors also thank Mr. Xiao DI for generously providing the circular cylinder flow dataset used in this study.